Four charged particles are at the corners of a square of side a. Determine (a) the electric field at the location of charge q and (b) the total electric force exerted on q.
1
Expert's answer
2020-10-21T09:54:30-0400

q is the charge

a is the (magnitude) which is then squared

for "A_q," "E=\frac{ke\times3q}{a^{2}}\times [cos(0), 0]" ,

"B_{q}, E= \frac{ke\times4q}{2a^{2}}\times[ cos45, sin45]"

"C_{q}, E=\frac{ke\times3q}{a^{2}}\times[0, sin90]"

i like using (i and j instead of x and y). Sum of the electric fields is

"E_{q} = \frac{ke\times q}{a^{2}}\times(3i + (4cos45)i + 4sin45j + 3j"

"\frac{ke\times q }{a^{2}}\times(5.828i + 5.828j)"

"\frac{ke\times q }{a^{2}}\times8.24"

"tan\theta = \frac{opposite}{adjacent}= \frac{5.828}{5.828}=1"

"tan^{-1}1=45^{0}"

(b) The total electric force on q is "\frac{ke\times q }{a^{2}}\times(8.28)" at an angle of "45^{o}"